OpenStudy (anonymous):

14. Which term in the expansion of (92-2/9)18 will have 912?

OpenStudy (anonymous):

We have to use Binomial Expansion which I have forgot now.. Can anybody know about Binomial Expansion...???

OpenStudy (cwrw238):

r+1 th term = nCr a ^(n-r) b^r thats for ( a + b)^n

OpenStudy (cwrw238):

so (r+1)th term = 18Cr 92^(18-r)(- (2/9)^18)

OpenStudy (anonymous):

@cwrw238 , there it will come r in place of 18 in power of b???

OpenStudy (cwrw238):

yes the last part is (-2/9)^r not (-2/9)^18

OpenStudy (anonymous):

\[(r+1)^{th}=^nC_r.a^{n−r}.b^r\] Here n= 18, a = 92 and b = 2/9 \[912=^{18}C_r.(92)^{18−r}.(\frac{−2}{9})^r\]

OpenStudy (cwrw238):

yes thats correct

OpenStudy (anonymous):

I think somewhat that the question is wrong...

OpenStudy (cwrw238):

yes i agree